# Hilbert-Guided Block-Sparse Local Attention

### v1

@@ A.1 Full Window Attention Results
At low resolutions such as 56×56 with a window size of 7×7, HWA fails to outperform WSA, primarily due to the limited sparsity under this configuration, which is insufficient to yield meaningful acceleration. This observation indicates that surpassing the performance of a kernel operating solely on full blocks requires a certain degree of sparsity—that is, a sufficiently high proportion of empty blocks must be skipped. Moreover, even at high resolutions, achieving ideal acceleration still depends on selecting appropriate window and block sizes.
@@
Table LABEL:tab:allwin100 reports results on an A100. The A100 setup also uses CUDA 12.6 and PyTorch 2.7.0, and the test parameters are identical to those on the RTX 3080. Compared with the 3080 results, memory usage and sparsity remain unchanged because the attention pattern does not change. Runtimes differ and are faster on the A100, as expected. The A100 also supports higher resolutions such as $256\times 256$, where HWA shows a more pronounced speedup.
@@
Table 7 presents end-to-end results on RTX 3080 and A100 for window-attention variants, combining the corresponding reshape and QKV projection under different input and window configurations. At $56\times 56$ resolution, HWA does not achieve good speedups, mainly because the sparsity is low and the attention part is not significantly accelerated. As the resolution increases and the window size is chosen appropriately, the speedup becomes substantial, especially on A100, where at I=128, W=16, the acceleration approaches $3\times$.
@@ A.2 Full Slide/Neighborhood Attention Results
Attention Forward Backward Sparsity
Time Memory Time Memory
$I=56\times 56,W=7\times 7$ SA 5.85 ms 622 MB 22.92 ms 1835 MB 0%
SA (Flex) 0.56 ms 13 MB 1.89 ms 50 MB 80.17%
HSA (Flex) 0.32 ms 13 MB 1.32 ms 50 MB 87.84%
NA2D (Flex) 0.56 ms 13 MB 1.96 ms 50 MB 79.84%
HNA (Flex) 0.31 ms 13 MB 1.28 ms 50 MB 87.84%
NA2D (NAT) 0.21 ms 13 MB 2.65 ms 75 MB -
HNA (NAT) 0.12 ms 13 MB 1.03 ms 75 MB -
$I=64\times 64,W=9\times 9$ SA 12.65 ms 1332 MB OOM OOM 0%
SA (Flex) 0.49 ms 17 MB 2.21 ms 65 MB 84.96%
HSA (Flex) 0.28 ms 17 MB 1.39 ms 65 MB 90.82%
NA2D (Flex) 0.49 ms 17 MB 2.16 ms 65 MB 84.38%
HNA (Flex) 0.28ms 17 MB 1.39 ms 65 MB 90.82%
NA2D (NAT) 0.28 ms 17 MB 3.52 ms 98 MB -
HNA (NAT) 0.20 ms 17 MB 1.37 ms 98 MB -
$I=96\times 96,W=9\times 9$ SA OOM OOM OOM OOM 0%
SA (Flex) 1.68 ms 37 MB 7.32 ms 146 MB 88.73%
HSA (Flex) 0.61 ms 37 MB 3.32 ms 146 MB 95.87%
NA2D (Flex) 1.71 ms 37 MB 7.53 ms 146 MB 88.58%
HNA (Flex) 0.61 ms 37 MB 3.10 ms 146 MB 95.87%
NA2D (NAT) 0.59 ms 37 MB 7.87 ms 219 MB -
HNA (NAT) 0.44 ms 37 MB 2.98 ms 219 MB -
$I=96\times 96,W=11\times 11$ SA OOM OOM OOM OOM 0%
SA (Flex) 1.83 ms 37 MB 7.82 ms 146 MB 87.89%
HSA (Flex) 0.61 ms 37 MB 3.23 ms 146 MB 95.87%
NA2D (Flex) 1.86 ms 37 MB 8.25 ms 146 MB 87.50%
HNA (Flex) 0.61ms 37 MB 3.19 ms 146 MB 95.87%
NA2D (NAT) 1.07 ms 37 MB 8.38 ms 219 MB -
HNA (NAT) 0.51 ms 37 MB 3.06 ms 219 MB -
$I=96\times 96,W=17\times 17$ SA OOM OOM OOM OOM 0%
SA (Flex) 2.79 ms 37 MB 11.79 ms 146 MB 81.06%
HSA (Flex) 0.98 ms 37 MB 4.50 ms 146 MB 93.17%
NA2D (Flex) 2.87 ms 37 MB 12.45 ms 146 MB 80.40%
HNA (Flex) 0.98 ms 37 MB 4.50 ms 146 MB 93.17%
NA2D (NAT) 1.22 ms 37 MB 9.08 ms 219 MB -
HNA (NAT) 0.77 ms 37 MB 4.87 ms 219 MB -
$I=128\times 128,W=17\times 17$ SA OOM OOM OOM OOM 0%
SA (Flex) 6.12 ms 66 MB 23.91 ms 260 MB 87.16%
HSA (Flex) 1.72 ms 66 MB 8.15 ms 260 MB 96.13%
NA2D (Flex) 6.41 ms 66 MB 26.06 ms 260 MB 86.72%
HNA (Flex) 1.72 ms 66 MB 8.17 ms 260 MB 96.13%
NA2D (NAT) 2.18 ms 66 MB 16.05 ms 390 MB -
HNA (NAT) 1.38 ms 66 MB 8.60 ms 390 MB -
@@
Table 9 reports all results for the Slide/Neighborhood Attention variants on A100. The same library versions and parameter settings used on the RTX 3080 are employed. The A100 can handle computations at $96\times 96$ resolution, but OOM occurs when the kernel size is further increased. Even so, HSA and HNA (Flex) still show strong speedups. Due to hardware resource constraints, the NATTEN library could not be installed on the A100 machine, so NAT-related attention variants were not evaluated.
@@ A.3 The Influence of Other Parameters
Figure 8 illustrates the impact of batch size, number of heads, and head dimension on speed. These factors do not alter sparsity, but they do impact runtime, and the optimal settings are hardware-dependent. On RTX 3080, when the batch size is 1, WSA and HSA run at nearly the same speed. As the batch size increases, the speedup grows, but it stops improving beyond 32, likely due to memory and bandwidth limits. The number of heads has little impact on speedup, whereas increasing the head dimension slows computation because it is more bandwidth-bound.
@@ A.4 SDPA Comparison
Figure 9: SDPA vs. HWA on RTX 3080.
Figure 9 compares SDPA (F.scaled_dot_product_attention)(F.scaled\_dot\_product\_attention) in FlexAttention with HWA across different resolutions and window sizes. SDPA can automatically dispatch to FlashAttention or memory-efficient kernels, accelerating dense attention such as WSA. When sparsity in FlexAttention is high enough, HWA remains faster than SDPA. At high resolutions, SDPA benefits from larger windows and can reach speeds comparable to HWA. For SA/NA, SDPA still materializes intermediates and remains bandwidth and memory-limited, so HSA/HNA perform better.

### v2

@@ A.1 Full Window Attention Results
At low resolutions such as 56×56 with a window size of 7×7, HWA (Flex) fails to outperform WSA, primarily due to the limited sparsity under this configuration, which is insufficient to yield meaningful acceleration. This observation indicates that surpassing the performance of a kernel operating solely on full blocks requires a certain degree of sparsity—that is, a sufficiently high proportion of empty blocks must be skipped. Moreover, even at high resolutions, achieving ideal acceleration still depends on selecting appropriate window and block sizes.
@@
Table 7 reports results on an A100. The A100 setup also uses CUDA 12.6 and PyTorch 2.7.0, and the test parameters are identical to those on the RTX 3080. Compared with the 3080 results, memory usage and sparsity remain unchanged because the attention pattern does not change. Runtimes differ and are faster on the A100, as expected. The A100 also supports higher resolutions such as $256\times 256$, where HWA (Flex) shows a more pronounced speedup.
@@
Table 8 presents end-to-end results on RTX 3080 and A100 for window-attention variants, combining the corresponding reshape and QKV projection under different input and window configurations. At $56\times 56$ resolution, HWA (Flex) does not achieve good speedups, mainly because the sparsity is low and the attention part is not significantly accelerated. As the resolution increases and the window size is chosen appropriately, the speedup becomes substantial, especially on A100, where at I=128, W=16, the acceleration approaches $3\times$.
@@ A.2 Full Slide/Neighborhood Attention Results
Table 9 reports all results for the Slide/Neighborhood Attention variants on A100. The same library versions and parameter settings used on the RTX 3080 are employed. The A100 can handle computations at $96\times 96$ resolution, but OOM occurs when the kernel size is further increased. Even so, HSA (Flex) and HNA (Flex) still show strong speedups. Due to hardware resource constraints, the NATTEN library could not be installed on the A100 machine, so NAT-related attention variants were not evaluated.
@@
Attention Forward Backward Sparsity
Time Memory Time Memory
$I=56\times 56,W=7\times 7$ SA 5.85 ms 622 MB 22.92 ms 1835 MB 0%
HSA (FA2) 0.25 ms 13 MB 0.75 ms 76 MB -
SA (xFormers) 1.31 ms 141 MB 3.56 ms 208 MB -
HSA (xFormers) 0.28 ms 13 MB 0.79 ms 76 MB -
SA (Flex) 0.56 ms 13 MB 1.89 ms 50 MB 80.17%
HSA (Flex) 0.32 ms 13 MB 1.32 ms 50 MB 87.84%
NA2D (Flex) 0.56 ms 13 MB 1.96 ms 50 MB 79.84%
HNA (Flex) 0.31 ms 13 MB 1.28 ms 50 MB 87.84%
NA2D (NAT) 0.21 ms 13 MB 2.65 ms 75 MB -
HNA (NAT) 0.12 ms 13 MB 1.03 ms 75 MB -
$I=64\times 64,W=9\times 9$ SA 12.65 ms 1332 MB OOM OOM 0%
HSA (FA2) 0.31 ms 17 MB 0.95 ms 97 MB -
SA (xFormers) 2.47 ms 204 MB 6.00 ms 306 MB -
HSA (xFormers) 0.37 ms 17 MB 1.02 ms 98 MB -
SA (Flex) 0.49 ms 17 MB 2.21 ms 65 MB 84.96%
HSA (Flex) 0.28 ms 17 MB 1.39 ms 65 MB 90.82%
NA2D (Flex) 0.49 ms 17 MB 2.16 ms 65 MB 84.38%
HNA (Flex) 0.28ms 17 MB 1.39 ms 65 MB 90.82%
NA2D (NAT) 0.28 ms 17 MB 3.52 ms 98 MB -
HNA (NAT) 0.20 ms 17 MB 1.37 ms 98 MB -
$I=96\times 96,W=9\times 9$ SA OOM OOM OOM OOM 0%
HSA (FA2) 0.69 ms 37 MB 2.08 ms 218 MB -
SA (xFormers) 4.79 ms 462 MB 13.49 ms 693 MB -
HSA (xFormers) 0.79 ms 37 MB 2.22 ms 219 MB -
SA (Flex) 1.68 ms 37 MB 7.32 ms 146 MB 88.73%
HSA (Flex) 0.61 ms 37 MB 3.32 ms 146 MB 95.87%
NA2D (Flex) 1.71 ms 37 MB 7.53 ms 146 MB 88.58%
HNA (Flex) 0.61 ms 37 MB 3.10 ms 146 MB 95.87%
NA2D (NAT) 0.59 ms 37 MB 7.87 ms 219 MB -
HNA (NAT) 0.44 ms 37 MB 2.98 ms 219 MB -
$I=96\times 96,W=11\times 11$ SA OOM OOM OOM OOM 0%
HSA (FA2) 0.69 ms 37 MB 2.08 ms 218 MB -
SA (xFormers) 6.58 ms 545 MB 18.71 ms 818 MB -
HSA (xFormers) 0.81 ms 37 MB 2.22 ms 219 MB -
SA (Flex) 1.83 ms 37 MB 7.82 ms 146 MB 87.89%
HSA (Flex) 0.61 ms 37 MB 3.23 ms 146 MB 95.87%
NA2D (Flex) 1.86 ms 37 MB 8.25 ms 146 MB 87.50%
HNA (Flex) 0.61ms 37 MB 3.19 ms 146 MB 95.87%
NA2D (NAT) 1.07 ms 37 MB 8.38 ms 219 MB -
HNA (NAT) 0.51 ms 37 MB 3.06 ms 219 MB -
$I=96\times 96,W=17\times 17$ SA OOM OOM OOM OOM 0%
HSA (FA2) 1.04 ms 37 MB 3.21 ms 218 MB -
SA (xFormers) 14.65 ms 884 MB 43 ms 1326 MB -
HSA (xFormers) 1.11 ms 37 MB 3.36 ms 219 MB -
SA (Flex) 2.79 ms 37 MB 11.79 ms 146 MB 81.06%
HSA (Flex) 0.98 ms 37 MB 4.50 ms 146 MB 93.17%
NA2D (Flex) 2.87 ms 37 MB 12.45 ms 146 MB 80.40%
HNA (Flex) 0.98 ms 37 MB 4.50 ms 146 MB 93.17%
NA2D (NAT) 1.22 ms 37 MB 9.08 ms 219 MB -
HNA (NAT) 0.77 ms 37 MB 4.87 ms 219 MB -
$I=128\times 128,W=17\times 17$ SA OOM OOM OOM OOM 0%
HSA (FA2) 1.86 ms 66 MB 5.68 ms 388 MB -
SA (xFormers) OOM OOM OOM OOM -
HSA (xFormers) 1.99 ms 66 MB 5.91 ms 390 MB -
SA (Flex) 6.12 ms 66 MB 23.91 ms 260 MB 87.16%
HSA (Flex) 1.72 ms 66 MB 8.15 ms 260 MB 96.13%
NA2D (Flex) 6.41 ms 66 MB 26.06 ms 260 MB 86.72%
HNA (Flex) 1.72 ms 66 MB 8.17 ms 260 MB 96.13%
NA2D (NAT) 2.18 ms 66 MB 16.05 ms 390 MB -
HNA (NAT) 1.38 ms 66 MB 8.60 ms 390 MB -
@@ A.3 The Influence of Other Parameters
Figure 8 illustrates the impact of batch size, number of heads, and head dimension on speed. These factors do not alter sparsity, but they do impact runtime, and the optimal settings are hardware-dependent. On RTX 3080, when the batch size is 1, WSA and HSA (Flex) run at nearly the same speed. As the batch size increases, the speedup grows, but it stops improving beyond 32, likely due to memory and bandwidth limits. The number of heads has little impact on speedup, whereas increasing the head dimension slows computation because it is more bandwidth-bound.
@@ A.4 SDPA Comparison
Figure 9: SDPA vs. HWA (Flex) on RTX 3080.
Figure 9 compares HWA (Flex) in FlexAttention with SDPA (F.scaled_dot_product_attention)(F.scaled\_dot\_product\_attention) across different resolutions and window sizes. SDPA can automatically dispatch to FlashAttention or memory-efficient kernels, accelerating dense attention such as WSA. When sparsity in FlexAttention is high enough, HWA (Flex) remains faster than SDPA. At high resolutions, SDPA benefits from larger windows and can reach speeds comparable to HWA (Flex). For SA/NA, SDPA still materializes intermediates and remains bandwidth and memory-limited, so HSA/HNA perform better.
A.5 Code
The code is open-sourced and available at https://github.com/Yunge6666/Hilbert-Local-Attention.
A.6 The Use of Large Language Models
This paper uses LLM for grammar correction and polishing.